Presented below is information related to Monty Company.

1. On July 6, Monty Company acquired the plant assets of Doonesbury Company, which had discontinued operations. The appraised value of the property is: Land $300,000 Buildings 900,000 Equipment 600,000 Total $1,800,000 Monty Company gave 12,300 shares of its $100 par value common stock in exchange. The stock had a market price of $168 per share on the date of the purchase of the property.

2. Monty Company expended the following amounts in cash between July 6 and December 15, the date when it first occupied the building. (Prepare consolidated entry for all transactions below.) Repairs to building $126,000 Construction of bases for equipment to be installed later 162,000 Driveways and parking lots 146,400 Remodeling of office space in building, including new partitions and walls 193,200 Special assessment by city on land 21,600

3. On December 20, the company paid cash for equipment, $312,000, subject to a 2% cash discount, and freight on equipment of $12,600.

Prepare entries on the books of Monty Company for these transactions. (Round intermediate calculations to 5 decimal places, e.g. 1.25124 and final answer to 0 decimal places e.g. 58,971. Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)
Answer #1

Date

Particulars

Debit ($)

Credit ($)

1.

Land

344,400

Building

1033,200

Equipment

688,800

       Common stock

1230,000

       Paid in capital in excess of par

836,400

(To record acquisition of plant assets by equity)

2.

Buildings

319,200

Equipment

162,000

Land Improvements

146,400

Land

21,600

     Cash

649,200

(To record the capital expenditures)

3

Equipment

318,360

      Cash

318,360

(To record purchase of equipment)

Working Notes

1.

Market Value of Stock: 12300 x $168 = $2,066,400

Common Stock: 12300 x 100 = $1,230,000

Paid in Capital in Excess of Common Stock: 2,066,400 –1,230,000 = $ 836,400

Amounts to include on statement –

Land: $2,066,400 X 300,000/1800,000 = $ 344,400

Buildings: $2,066,400 X 900,000/1800,000 = $ 1,033,200

Equipment: $2,066,400 X 600,000/1800,000 = $ 688,800

2.

Buildings = repairs + remodeling of office space = 126,000 + 193,200 = $ 319,200

Construction of bases for equipment is to be capitalized to equipment

Driveways and parking lots is to be capitalized to land improvements

Special assessment by city on land is to be added to cost of land.

3.

Cash discount is to be reduced from the purchase price of equipment.

Further, freight is to be added to the cost of equipment as it aids in bringing equipment to the condition for use, simply, being freight inward for assets.

Hence, equipment = 312,000 * 98 % +12,600 = $ 318,360

kindly upvote